Let F(x)=∫

t−3

t

2+7

for − ∞ < x < ∞

x


(a) Find the value of x where F attains its minimum value.

(b) Find intervals over which F is only increasing or only decreasing.

(c) Find open intervals over which F is only concave up or only concave down.



1
Expert's answer
2022-01-04T08:05:24-0500
"F(x)=\displaystyle\int_{0}^x(t-3t^2+7)dt"


By the Fundamental Theorem of Calculus


"F'(x)=x-3x^2+7"

Find the critical points of "F(x)"

"F'(x)=0=>x-3x^2+7=0"

"3x^2-x-7=0"

"D=(-1)^2-4(3)(-7)=85"

"x=\dfrac{1\pm\sqrt{85}}{2(3)}"

"x_1=\dfrac{1-\sqrt{85}}{6}, x_2=\dfrac{1+\sqrt{85}}{6}"

If "x<\dfrac{1-\sqrt{85}}{6}, F'(x)<0, F(x)" decreases.

If "\dfrac{1-\sqrt{85}}{6}<x<\dfrac{1+\sqrt{85}}{6}, F'(x)>0, F(x)" increases.

If "x>\dfrac{1+\sqrt{85}}{6}, F'(x)<0, F(x)" decreases.


(a)The function "F(x)" has the local minimum at "x=\dfrac{1-\sqrt{85}}{6}."


(b) The function "F(x)" increases on

"(\dfrac{1-\sqrt{85}}{6}, \dfrac{1+\sqrt{85}}{6})."

The function "F(x)" decreases on

"(-\infin, \dfrac{1-\sqrt{85}}{6})\cup(\dfrac{1+\sqrt{85}}{6}, \infin)"

(c)


"F''(x)=(x-3x^2+7)'=1-6x"

Find the inflection point(s)


"F''(x)=0=>1-6x=0=>x=\dfrac{1}{6}"

If "x<\dfrac{1}{6}, F''(x)>0, F(x)" is concave up.

If "x>\dfrac{1}{6}, F''(x)<0, F(x)" is concave down.

The function "F(x)" is only concave up on "(-\infin, \dfrac{1}{6})."

The function "F(x)" is only concave down on "(\dfrac{1}{6}, \infin)."